OpenStudy (anonymous):

The length of a rectangle is 3 more than the width. The perimeter of the rectangle is 44. Find the length and the width.

OpenStudy (anonymous):

you take x as the width so then the length is 3x. then (3x*2)+(x*2)=44 and you solve for x

OpenStudy (anonymous):

thank

OpenStudy (anonymous):

still can't solve it

OpenStudy (anonymous):

well 3x*2 is 6x and x*2 is 2x so 6x+2x=44, 8x=44,x=44/8

OpenStudy (anonymous):

and then length is 3x so length=44/8*3

OpenStudy (anonymous):

3x can't time 2

OpenStudy (anonymous):

2 don't have x behind

OpenStudy (anonymous):

yes 3x multiplied by 2 is 6x

OpenStudy (anonymous):

 perimeter = 2(length + breadth) 44=2(l+b) 22=l+b length=3+x breadth = x therefore 22=3+x+x 22=3+2x 22-3=2x 19=2x 19/2=x(width) (19/2)+3=length

OpenStudy (anonymous):

i still don't get it

OpenStudy (anonymous):

u got it now?